OpenStudy (anonymous):

Please help! Graph a hyperbola with foci at (0, ±10) and vertices at (0, ±8) using graphing technology.

OpenStudy (amistre64):

i assume the graphing technology needs an equation for the hyperB?

OpenStudy (amistre64):

the general setup for a hyperB is: \[\frac{(x-c_x)^2}{a^2}-\frac{(y-c_y)^2}{b^2}=1\] or \[\frac{(y-c_y)^2}{a^2}-\frac{(x-c_x)^2}{b^2}=1\] depending on how the hyperB opens

OpenStudy (anonymous):

yes, but i don't understand how i find those variables with what the question gives me /:

OpenStudy (amistre64):

well, half the distance between your vertexes is "a" the midpoint between your vertexis gives you (Cx,Cy) your centers for starters

OpenStudy (amistre64):

by convention; "c" is the measure from center to focus; such that: a^2 + b^2 = c^2 if i recall it right

OpenStudy (amistre64):

can you tell me what you thing the center point for this thing would be?

OpenStudy (anonymous):

but how do i find a and b without an equation?

OpenStudy (amistre64):

you are given the values for a and c; and use the pythag thrm to find b

OpenStudy (anonymous):

like i would have a^2 = 144 a = 12 ... and lets say i was finding a minor axis, i would do 2a ... 2(12) = 24

OpenStudy (amistre64):

jsut remember that "a" is the distance from center to vertex; you are given both vertex points to begin with .... right?

OpenStudy (anonymous):

(0, 8) and (0, -8) are the verices , yes

OpenStudy (amistre64):